OpenStudy (anonymous):

Can you help me solve? The circus is coming to town. Sue is selling tickets. The first day she sold 4 senior citizen tickets and 2 children's tickets for a total of $80. The second day she received $145 for 5 senior citizen tickets and 7 children's tickets. What is the price of a child's ticket and a senior citizen ticket? I know that 4s + 2c = 80 and 5s + 7c = 145, but I don't know where to go from here. I want to know how to fully solve it

OpenStudy (anonymous):

4s+2c=80 2s+c=40 c=40-2s put c value in 2nd eq. 5s+7(40-2s)=145 5s+280-14s=145 -9s=145-280 -9s=-135 9s=135 s=15 and c=40-2.15 c=40-30 c=10

OpenStudy (anonymous):

how did you get 2x+c=40 for the second equation?

OpenStudy (anonymous):

@Probhat

OpenStudy (anonymous):

4s+2c=80 (2s+c)=80/2 2s+c=40....it's not the second eq. second one is given below line 5.

OpenStudy (anonymous):

@haileemackk

OpenStudy (anonymous):

oh okay, sorry for the confusion! thank you!